# IMPACT[Uncaptioned image]: Intelligent Motion Planning with Acceptable Contact Trajectories via Vision-Language Models

### v2

@@ I Introduction
Classical motion planning for robot manipulation [25, 37] frames the problem as finding a path for the robot’s end-effector to reach a target while avoiding collisions with obstacles. This formulation is generally desirable, but can be highly restrictive, especially in densely cluttered environments [7]. In such cases, some incidental contact may be necessary to achieve a task, or to accomplish it more efficiently than by strictly avoiding all collisions.
@@
To address this challenge, we propose Intelligent Motion Planning with Acceptable Contact Trajectories (IMPACT), a novel framework that leverages modern Vision-Language Models (VLMs) such as GPT-4o [29] to infer initial semantic costs for objects in the scene. However, knowing that an object is generally safe to touch is different from knowing how and where to push it without causing a cascade of collisions. Our method repeatedly samples push outcomes to estimate the probability of causing a collision with obstacles or the target. This risk is condensed into safety scores that are used to construct an anisotropic cost map, where interaction costs vary with direction since pushing an object from one side versus another can lead to different outcomes. We integrate this directional cost map with a contact-aware A* planner [11] that computes paths to reach targets. We perform experiments on contact-rich, densely cluttered reaching tasks in simulation and real-world settings. We evaluate IMPACT using multiple quantitative metrics, including human preference rankings from a user study. Results indicate that users prefer IMPACT over alternatives, suggesting that it is a promising approach for facilitating semantically-acceptable contact-rich manipulation.
@@ II-A Robot Motion Planning
Classical motion planning algorithms can be broadly characterized as optimization-based or sampling-based [37]. Optimization-based methods, such as TrajOpt [34], frame the problem as minimizing a cost function subject to kinematic, dynamics, and obstacle avoidance constraints. Sampling-based methods, such as PRM [15], RRT [20], and RRT* [14], incrementally build a graph or tree of feasible paths. While there are numerous variants of these methods, a common underlying theme is the constraint of avoiding any collisions. This is generally desirable, but has limitations. In densely cluttered scenarios where contact is inevitable, these motion planners might not find a solution, even though one might exist if the robot engages in light contact. Our method seamlessly integrates with prior motion planning algorithms for flexible contact-rich manipulation.
Some motion planning works modify cost functions so that certain obstacles, such as leaves in foliage, are permeable [17, 28] and thus allow for some contact. These methods have been applied to domain-specific foliage settings and not tested in common cluttered scenarios with rigid objects. In closely-related work, Xie et al. [41] propose to use semantic language commands to enable semantically-acceptable contact. Unlike [41], we do not require explicit language instructions about which contacts are acceptable or not, since we utilize VLMs to automate this process.
@@ II-B Contact-Rich Robot Manipulation
Dealing with contacts is challenging in manipulation [37]. This may refer to frequent contact between an object (that the robot grips/holds) and an environment, such as tight placement tasks like connector and peg insertion as explored in robotic reinforcement learning works [22, 33, 21]. Another method for contact-rich problems is extrinsic dexterity [4, 42], which takes advantage of contacts between an object and rigid parts of the environment (such as walls) to reorient the object to improve subsequent manipulation. Other works consider contact-rich interactions in different applications such as assistive robots that incorporate human feedback and touch [13]. In contrast, we consider the relatively less-explored “contact-rich” setting, where “contact” refers to robot parts touching the environment’s obstacles.
@@ II-C Vision-Language Models (VLMs) for Robotics
VLMs such as GPT-4o [29] and Gemini [10] are trained on broad Internet-scale data and have remarkable semantic and spatial knowledge. Thus, the research community has explored numerous applications of VLMs in robotics [9, 16]. One way to use VLMs in robotics is to generate high-level task plans, in the form of natural language [1, 8] or executable code [23, 35]. Another way to use VLMs is for reward [26, 36] or task design [38, 39]. VLMs can also aid low-level affordance reasoning [18, 24] which, in an extension of this direction, can involve generating full trajectories [19]. In contrast to these works, we use VLMs for a complementary objective: to infer object contact tolerances to guide contact-rich motion planning.
In closely-related work, VoxPoser [12] exploits VLMs for open-world reasoning and visual grounding to compose a 3D value map to guide robotic interactions. As in [12], we use VLMs to assign semantic costs to objects. However, unlike VoxPoser, which requires the user to explicitly specify which objects to avoid with language, e.g., “get the item, but watch out for that vase!” we do not require explicit language commands. Instead, we leverage the improving spatial and semantic knowledge of recent VLMs [5] to determine object contact tolerances, and show manipulation in more densely cluttered environments. Other recent work relies on VLMs for semantically-safe manipulation [2] but assumes that any collisions are undesirable.
@@ IV Method: IMPACT
Figure 2: Overview of IMPACT. There is a toy bear, a coffee cup and a glue bottle (target) on the table. The VLM receives an annotated image $I^{\prime}$ and a language template prompt $\ell$ with object information from SAM2 [31], and outputs costs for the three objects. We use a cost of 1-1 for the target object. We construct a 3D voxel grid $V$ using these costs and then flatten it to produce an anisotropic, contact-aware cost map $M^{\prime}$. The contact-aware A* planner searches over three motion primitives in this map: Move, Push and Rotate to generate a trajectory. The planner’s state space includes the robot’s end-effector pose and the displaced positions of low-cost objects. These guide the robot to avoid the coffee cup but make contact with the toy bear at the appropriate direction to reach the glue bottle.
@@ IV-A Obtaining Object Costs using a VLM
We use GPT-4o (hereafter, GPT) as the VLM to generate the cost of all objects in the scene due to its strong spatial reasoning capabilities [5], but our approach is compatible with other VLMs. The input to the VLM includes both an annotated image and a text prompt. We mount an RGBD camera to capture a scene image $I$, and use SAM2 [31] to segment and annotate the objects. The segmented and annotated image $I^{\prime}$ is ultimately provided as part of the input to GPT. We also design a text prompt template $\ell$ which includes the list of objects in the scene (but is otherwise task-agnostic), and some general principles related to the concept of contact tolerance. Objects are labeled with the same numbers as in the input image, marked by SAM2. The full text prompt can be found on our website. The output of GPT is a dictionary with the cost of all queried objects. Fig. 1 (second row) shows an example of object costs generated from GPT. It assigns a high cost of $8$ to the wine glass and a lower cost of $3$ to the toy bear, which indicates different tolerances to contact across objects.
@@ IV-B2 Contact-Aware A* Planning
To generate a plan, we search for a sequence of primitives that guides the end-effector to the target while accounting for directional costs in $M^{\prime}$. We formulate this as a graph search problem and use A* [11], which is well-suited for this setting and provides deterministic control compared to sampling-based algorithms. The planner’s state is defined as $S=(p,r,\mathcal{D})$, where $(p,r)$ is the robot end-effector position and orientation in 2D space and $\mathcal{D}$ is a list that tracks the cumulative 2D displacement of low-cost objects. At each step, the planner expands nodes using the three primitives. For Push, we estimate the object displacement using $d=D\cdot\big(\cos(\varphi)\big)^{\gamma}$, where $D$ is the maximum displacement when pushing in the optimal direction, $\varphi$ is the angular deviation between the push direction and the object’s reverse surface normal at the contact point, and the decay parameter $\gamma$ is set to $1.3$. IMPACT extends standard A* search by incorporating world state changes caused by pushes. Each valid Push creates a new branch in the search tree where an object has been permanently displaced to a new position. The resulting state, $S^{\prime}=(p^{\prime},r^{\prime},\mathcal{D^{\prime}})$, contains the updated robot pose and an updated world configuration.
@@
where $C(a)$ is the action cost and $P(s^{\prime})$ is a gripper placement penalty. We define action costs based on the primitive. Move assigns a cost proportional to the Euclidean distance between two end-effector positions. Rotate applies a fixed cost for orientation changes. Push generates its cost from the value $M^{\prime}[x,y]$ at the contact point $(x,y)$. The penalty $P(s^{\prime})$ consists of three terms: (i) the average cost of the region the gripper overlaps, (ii) a collision penalty if the gripper collides with an object, and (iii) a proximity penalty proportional to the distance to high-cost objects. We defer additional details to the Appendix. By combining these costs, the planner can find a trajectory that minimizes risk while permitting acceptable contact when necessary; see Fig. 3 for more information about the A* planner. The path is then converted back to 3D space for evaluation and execution.
@@ V-A Experiment Setup in Simulation
We build and test our pipeline using PyBullet simulation [6]. In simulation, we create 20 scenes designed to test contact-rich manipulation. These use a hybrid object dataset, which includes tall and bulky items (e.g., sugar boxes and water pitchers) from the YCB [3] dataset and fragile objects (e.g., wine glasses and stacked bowls) using 3D models generated from TRELLIS [40]. TRELLIS allows constructing delicate object meshes from single-view images and introduces fragility constraints not present in YCB objects. All objects are placed on a tabletop.
@@
TABLE I: Comparison of path planning algorithms and results in PyBullet simulation. We report five quantitative metrics (see Sec. V-D). “Success Rate” only counts the trajectories that reach the target and satisfy thresholds on “Path Cost,” “Contact Duration,” and “Unsafe Object (human-designated) Displacement.” The arrow $\uparrow$ indicates larger values of the metric correspond to better performance, and $\downarrow$ represents the opposite. Collision-free baselines and LAPP [41] do not use VLMs to generate object costs, so they do not have values for “Path Cost” in this table.
@@ V-B Motion Planning Methods
As presented in Sec. IV-B, IMPACT performs directional contact-aware path planning on the flattened cost map $M^{\prime}$. In simulation, we also test the 3D cost map $V$ with RRT [20] and RRT* [14] as the motion planners, and thus call the respective methods VLM Cost+RRT and VLM Cost+RRT*. Each trajectory consists of a sequence of robot end-effector positions. The total cost of a trajectory $\tau$ is the sum of costs of obstacles the robot arm collides with, $Cost(\tau)=\sum_{i=1}^{|\tau|}V[\tau_{i}]$}. Here, $|\tau|$ denotes the length of the trajectory $\tau$, and $\tau_{i}$ is the position of the $i$-th waypoint. In addition, we test another planning method that directly applies A* search without any directional push safety analysis. We refer to it as VLM Cost+A* w/o direction. It only relies on the isotropic costs in the flattened cost map $M$.
@@ V-C2 Language-Conditioned Path Planning (LAPP)
We use LAPP [41] as a strong baseline, because (like IMPACT) it also allows robots to make collisions with specific objects in the environment. LAPP trains a language-conditioned collision function that predicts whether a robot will collide with objects other than the one specified in a language instruction (which the robot is permitted to collide with). The collision function has three inputs: (i) CLIP [30] image embeddings of the scene, (ii) CLIP text embeddings of the language instruction (e.g., “can collide with toys”), and (iii) joint configurations of the robot arm. We build directly upon the official open-source LAPP code.
@@ V-D Evaluation Metrics
Figure 5: Examples of trajectories planned by IMPACT (top row) and LAPP (bottom row) in PyBullet simulation [6]. The obstacles are: a coke bottle, a pitcher, a sugar box and a pile of bowls. The target object is the mug behind the obstacles. The planned paths are shown in an overlaid green curve in each image. We also provide LAPP with a language instruction “Can collide with the pitcher and the sugar box.” See Sec. V-E for more details.
@@ V-D1 User Study Evaluation
To better assess IMPACT’s ability to generate semantically acceptable behavior, we conduct a user study approved by the IRB office of our institution. We develop a website to collect human feedback, where participants evaluate robot trajectories by watching videos (see Fig. 4). Each question presents two videos of robot trajectories in the same scene with the same target object. One video is generated by an algorithm that utilizes VLM-assigned costs, and the other by the corresponding Collision-Free baseline with the same planner or by LAPP. Users are not informed which algorithm produced each trajectory. Following a similar approach as Mirjalili et al. [27], participants are asked to select the trajectory they find most acceptable.
@@ VI-A Experiment Setup
We mount two Intel RealSense D435 cameras on both sides of the scene to capture RGBD images and generate voxel grids. In the real world, we need to get each object’s position to assign the associated object cost to those voxels. Unlike in simulation, we use Grounded SAM 2 [31, 32] to generate the segmented point cloud of each object as we build the cost map. Each image and the object list are provided as input to Grounded SAM 2 to predict the segmentation masks of the objects. We then get a multi-view segmented point cloud and convert it to the cost map. The motion planning part is the same as in simulation.
For physical experiments, we test IMPACT and LAPP. To adapt LAPP for real-world experiments, we fine-tune it on a small dataset containing 25 real-world RGB images, manually annotated with collision scores. The images are captured by an Intel RealSense D435 camera that faces both the robot and the scene. The annotations explicitly label which objects are safe to collide with (for example, “can collide with plastic bottle”). Objects not mentioned in the language prompts are treated as unsafe by default, aligning with LAPP’s methodology of explicit collision rules. Human operators pair each scene with joint configurations and task-aligned prompts, reflecting the original work’s use of free-form language constraints [41]. This retains LAPP’s pre-trained reasoning while grounding predictions in real-world spatial relationships and safety priorities. Of the 10 test scenes, 8 use objects seen during fine-tuning, while 2 scene contains novel objects to evaluate generalization.

### v3

@@ I Introduction
Classical motion planning for robot manipulation [undefv, undefai] frames the problem as finding a path for the robot’s end-effector to reach a target while avoiding collisions with obstacles. This formulation is generally desirable, but can be highly restrictive, especially in densely cluttered environments [undeff]. In such cases, some incidental contact may be necessary to achieve a task, or to accomplish it more efficiently than by strictly avoiding all collisions.
@@
To address this challenge, we propose Intelligent Motion Planning with Acceptable Contact Trajectories (IMPACT), a novel framework that leverages modern Vision-Language Models (VLMs) such as GPT-4o [undefz] to infer initial semantic costs for objects in the scene. However, knowing that an object is generally safe to touch is different from knowing how and where to push it without causing a cascade of collisions. Our method repeatedly samples push outcomes to estimate the probability of causing a collision with obstacles or the target. This risk is condensed into safety scores that are used to construct an anisotropic cost map, where interaction costs vary with direction since pushing an object from one side versus another can lead to different outcomes. We integrate this directional cost map with a contact-aware A* planner [undefi] that computes paths to reach targets. We perform experiments on contact-rich, densely cluttered reaching tasks in simulation and real-world settings. We evaluate IMPACT using multiple quantitative metrics, including human preference rankings from a user study. Results indicate that users prefer IMPACT over alternatives, suggesting that it is a promising approach for facilitating semantically-acceptable contact-rich manipulation.
@@ II-A Robot Motion Planning
Classical motion planning algorithms can be broadly characterized [undefai] as optimization-based such as TrajOpt [undefae], or sampling-based such as PRM [undefn], RRT [undefq], and RRT* [undefm]. While there are numerous variants of these methods, a common underlying theme is the constraint of avoiding any collisions. In densely cluttered scenarios where contact is inevitable, these motion planners might not find a solution, even though one might exist if the robot engages in light contact. Prior work like Navigation Among Movable Obstacles [undefah] and Minimum Constraint Removal [undefj, undefaj] reason about relocating or removing obstacles to obtain feasible paths. Our method seamlessly integrates with prior motion planning algorithms for flexible contact-rich manipulation, without requiring explicit object relocation or environment modification.
Some motion planning works modify cost functions so that certain obstacles, such as leaves in foliage, are permeable [undefo, undefy] and thus allow for some contact. These methods have been applied to domain-specific foliage settings and not tested in common cluttered scenarios with rigid objects. In closely-related work, Xie et al. [undefan] propose to use semantic language commands to enable semantically-acceptable contact. Unlike [undefan], we do not require explicit language instructions about which contacts are acceptable or not, since we utilize VLMs to automate this process.
@@ II-B Contact-Rich Robot Manipulation
Dealing with contacts is challenging in manipulation [undefai]. This may refer to frequent contact between an object (that the robot grips/holds) and an environment, such as tight placement tasks like connector and peg insertion as explored in robotic reinforcement learning works [undefs, undefad, undefr]. Another method for contact-rich problems is extrinsic dexterity [undefc, undefao], which takes advantage of contacts between an object and rigid parts of the environment (such as walls) to reorient the object to improve subsequent manipulation. Other works consider contact-rich interactions in different applications such as assistive robots that incorporate human feedback and touch [undefl]. In contrast, we consider the relatively less-explored “contact-rich” setting, where “contact” refers to robot parts touching the environment’s obstacles.
@@ II-C Vision-Language Models (VLMs) for Robotics
VLMs such as GPT-4o [undefz] and Gemini [undefh] are increasingly applied in robotics for high-level planning [undef, undefg, undeft, undefaf], reward or task design  [undefw, undefag, undefak, undefal], and low-level affordance reasoning [undefp, undefu]. In contrast to these works, we use VLMs specifically to infer object contact tolerances to guide contact-rich motion planning.
In closely-related work, VoxPoser [undefk] exploits VLMs for open-world reasoning and visual grounding to compose a 3D value map to guide robotic interactions. As in [undefk], we use VLMs to assign semantic costs to objects. However, unlike VoxPoser, which requires the user to explicitly specify which objects to avoid with language, e.g., “get the item, but watch out for that vase!” we do not require explicit language commands. Instead, we leverage the improving spatial and semantic knowledge of recent VLMs [undefd] to determine object contact tolerances, and show manipulation in more densely cluttered environments. Other recent work relies on VLMs for semantically-safe manipulation [undefa] but assumes that any collisions are undesirable.
@@ IV Method: IMPACT
Figure 2: Overview of IMPACT. There is a toy bear, a coffee cup and a glue bottle (target) on the table. The VLM receives an annotated image $I^{\prime}$ and a language template prompt $\ell$ with object information from SAM2 [undefab], and outputs costs for the three objects. We use a cost of 1-1 for the target object. We construct a 3D voxel grid $V$ using these costs and then flatten it to produce an anisotropic, contact-aware cost map $M^{\prime}$. The contact-aware A* planner searches over three motion primitives in this map: Move, Push and Rotate to generate a trajectory. The planner’s state space includes the robot’s end-effector pose and the displaced positions of low-cost objects. These guide the robot to avoid the coffee cup but make contact with the toy bear at the appropriate direction to reach the glue bottle.
@@ IV-A Obtaining Object Costs using a VLM
We use GPT-4o (hereafter, GPT) as the VLM to generate the cost of all objects in the scene due to its strong spatial reasoning capabilities [undefd], but our approach is compatible with other VLMs. The VLM is used in a zero-shot inference setting without any fine-tuning. The input to the VLM includes both an annotated image and a text prompt. We mount an RGBD camera to capture a scene image $I$, and use SAM2 [undefab] to segment and annotate the objects. The segmented and annotated image $I^{\prime}$ is ultimately provided as part of the input to GPT. We also design a text prompt template $\ell$ which includes the list of objects in the scene (but is otherwise task-agnostic), and some general principles related to the concept of contact tolerance. Objects are labeled with the same numbers as in the input image, marked by SAM2. The full text prompt can be found on our website. The output of GPT is a dictionary with the cost of all queried objects. Fig. 1 (second row) shows an example of object costs generated from GPT. It assigns a high cost of $8$ to the wine glass and a lower cost of $3$ to the toy bear, which indicates different tolerances to contact across objects.
@@ IV-B2 Contact-Aware A* Planning
To generate a plan, we search for a sequence of primitives that guides the end-effector to the target while accounting for directional costs in $M^{\prime}$. We formulate this as a graph search problem and use A* [undefi], which is well-suited for this setting and provides deterministic control compared to sampling-based algorithms. The planner’s state is defined as $S=(p,r,\mathcal{D})$, where $(p,r)$ is the robot end-effector position and orientation in 2D space and $\mathcal{D}$ is a list that tracks the cumulative 2D displacement of low-cost objects. At each step, the planner expands nodes using the three primitives. For Push, we estimate the object displacement using $d=D\cdot\big(\cos(\varphi)\big)^{\gamma}$, where $D$ is the maximum displacement when pushing in the optimal direction, $\varphi$ is the angular deviation between the push direction and the object’s reverse surface normal at the contact point, and the decay parameter $\gamma$ is set to $1.3$. IMPACT extends standard A* search by incorporating world state changes caused by pushes. Each valid Push creates a new branch in the search tree where an object has been permanently displaced to a new position. The resulting state, $S^{\prime}=(p^{\prime},r^{\prime},\mathcal{D^{\prime}})$, contains the updated robot pose and an updated world configuration.
@@
where $C(a)$ is the action cost and $P(s^{\prime})$ is a gripper placement penalty. We define action costs based on the primitive. Move assigns a cost proportional to the Euclidean distance between two end-effector positions. Rotate applies a fixed cost for orientation changes. Push generates its cost from the value $M^{\prime}[x,y]$ at the contact point $(x,y)$. The penalty $P(s^{\prime})$ consists of three terms: (i) the average cost of the region the gripper overlaps, (ii) a collision penalty if the gripper collides with an object, and (iii) a proximity penalty proportional to the distance to high-cost objects. For additional details of push analysis and the planner, please refer to the appendix. By combining these costs, the planner can find a trajectory that minimizes risk while permitting acceptable contact when necessary; see Fig. 3 for more information about the A* planner. The path is then converted back to 3D space for evaluation and execution.
@@ V-A Experiment Setup in Simulation
We build and test our pipeline using PyBullet simulation [undefe]. In simulation, we create 20 scenes designed to test contact-rich manipulation. These use a hybrid object dataset, which includes tall and bulky items (e.g., sugar boxes and water pitchers) from the YCB [undefb] dataset and fragile objects (e.g., wine glasses and stacked bowls) using 3D models generated from TRELLIS [undefam]. TRELLIS allows constructing delicate object meshes from single-view images and introduces fragility constraints not present in YCB objects. All objects are placed on a tabletop.
@@
TABLE I: Comparison of path planning algorithms and results in PyBullet simulation. We report five quantitative metrics (see Sec. V-D). “Success Rate” only counts the trajectories that reach the target and satisfy thresholds on “Path Cost,” “Contact Duration,” and “Unsafe Object (human-designated) Displacement.” The arrow $\uparrow$ indicates larger values of the metric correspond to better performance, and $\downarrow$ represents the opposite. Collision-free baselines and LAPP [undefan] do not use VLMs to generate object costs, so they do not have values for “Path Cost” in this table.
@@ V-B Motion Planning Methods
As presented in Sec. IV-B, IMPACT performs directional contact-aware path planning on the flattened cost map $M^{\prime}$. In simulation, we also test the 3D cost map $V$ with RRT [undefq] and RRT* [undefm] as the motion planners, and thus call the respective methods VLM Cost+RRT and VLM Cost+RRT*. Each trajectory consists of a sequence of robot end-effector positions. The total cost of a trajectory $\tau$ is the sum of costs of obstacles the robot arm collides with, $Cost(\tau)=\sum_{i=1}^{|\tau|}V[\tau_{i}]$}. Here, $|\tau|$ denotes the length of the trajectory $\tau$, and $\tau_{i}$ is the position of the $i$-th waypoint. In addition, we test another planning method that directly applies A* search without any directional push safety analysis. We refer to it as VLM Cost+A* w/o direction. It only relies on the isotropic costs in the flattened cost map $M$.
@@ V-C2 Language-Conditioned Path Planning (LAPP)
We use LAPP [undefan] as a strong baseline, because (like IMPACT) it also allows robots to make collisions with specific objects in the environment. LAPP trains a language-conditioned collision function that predicts whether a robot will collide with objects other than the one specified in a language instruction (which the robot is permitted to collide with). The collision function has three inputs: (i) CLIP [undefaa] image embeddings of the scene, (ii) CLIP text embeddings of the language instruction (e.g., “can collide with toys”), and (iii) joint configurations of the robot arm. We build directly upon the official open-source LAPP code.
@@ V-D Evaluation Metrics
Figure 5: Examples of trajectories planned by IMPACT (top row) and LAPP (bottom row) in PyBullet simulation [undefe]. The obstacles are: a coke bottle, a pitcher, a sugar box and a pile of bowls. The target object is the mug behind the obstacles. The planned paths are shown in an overlaid green curve in each image. We also provide LAPP with a language instruction “Can collide with the pitcher and the sugar box.” See Sec. V-E for more details.
@@ V-D1 User Study Evaluation
To better assess IMPACT’s ability to generate semantically acceptable behavior, we conduct a user study approved by the IRB office of our institution. We develop a website to collect human feedback, where participants evaluate robot trajectories by watching videos (see Fig. 4). Each question presents two videos of robot trajectories in the same scene with the same target object. One video is generated by an algorithm that utilizes VLM-assigned costs, and the other by the corresponding Collision-Free baseline with the same planner or by LAPP. Users are not informed which algorithm produced each trajectory. Following a similar approach as Mirjalili et al. [undefx], participants are asked to select the trajectory they find most acceptable.
@@ VI-A Experiment Setup
We mount two Intel RealSense D435 cameras on both sides of the scene to capture RGBD images and generate voxel grids. In the real world, we need to get each object’s position to assign the associated object cost to those voxels. Unlike in simulation, we use Grounded SAM 2 [undefab, undefac] to generate the segmented point cloud of each object as we build the cost map. Each image and the object list are provided as input to Grounded SAM 2 to predict the segmentation masks of the objects. We then get a multi-view segmented point cloud and convert it to the cost map. The motion planning part is the same as in simulation.
For physical experiments, we test IMPACT and LAPP. To adapt LAPP for real-world experiments, we fine-tune it on a small dataset containing 25 real-world RGB images, manually annotated with collision scores. The images are captured by an Intel RealSense D435 camera that faces both the robot and the scene. The annotations explicitly label which objects are safe to collide with (for example, “can collide with plastic bottle”). Objects not mentioned in the language prompts are treated as unsafe by default, aligning with LAPP’s methodology of explicit collision rules. Human operators pair each scene with joint configurations and task-aligned prompts, reflecting the original work’s use of free-form language constraints [undefan]. This retains LAPP’s pre-trained reasoning while grounding predictions in real-world spatial relationships and safety priorities. Of the 10 test scenes, 8 use objects seen during fine-tuning, while 2 scene contains novel objects to evaluate generalization.